OpenStudy (anonymous):

@hoblos : Solve 6(x + 1) - 4 = 3x + 2.

OpenStudy (hoblos):

6(x + 1) - 4 = 3x + 2 6x + 6 -4 = 3x +2 6x +2 = 3x +2 3x =0 so x = 0